OpenStudy (anonymous):

i need help. 6 14 9 -- * -- * -- 7 27 10 mulitply. give answer too lowest term.

OpenStudy (anonymous):

6*14*9 = 756 7*27*10=1890 756/1890 = 2/5

OpenStudy (anonymous):

What part do you need help with